OpenStudy (anonymous):

The surface area of a cylinder is given by the formula SA = 2r2 + 2rh. A cylinder has a radius of 15 cm and a surface area of 1,770 cm2 . Find the height of the cylinder.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

sorry for asking you all the time

OpenStudy (igreen):

\(SA = 2r^2 + 2rh\) Is that 2 an exponent? the one after 2r.

OpenStudy (igreen):

It's fine, I'm on here to help people like you :P

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

its an exponent

OpenStudy (igreen):

\(SA = 2r^2 + 2rh\) Plug in what we know: \(1770 = 2(15)^2 + 2(15)h\) Simplify the exponent: \(1770 = 2(225) + 2(15)h\) Multiply: \(1770 = 450 + 2(15)h\) Multiply: \(1770 = 450 + 30h\) Subtract 450 to both sides: \(1320 = 30h\) Divide 30 to both sides: \(h = 44\)